# High Performance Distributed Control Law for Large-Scale Linear Systems: A Partitioned Distributed Observer Approach

### v3

@@ III-A Network partition
Algorithm 1 Partition algorithm of distributed observer for large-scale systems
0:  $\mathcal{N}_{i}$ and $\mathcal{C}_{i}$ $\forall i\in\mathcal{V}$
0:  A partition of distributed observer
1:  Establishing partitions based on greedy algorithm;
2:  Initialize $p=1$ and $\mathcal{O}_{p}=\emptyset$; initialize $\mathscr{P}_{i}=\emptyset$ $\forall i\in\mathcal{V}$;
3:  for $i=1$ to $N$ do
4:     Choose $r_{i}\in\mathcal{V}$;
5:     if $\mathcal{N}_{r_{i}}\backslash\mathcal{O}(\mathscr{P}_{r_{i}})\neq\emptyset$ then
6:        for $j\in\mathcal{N}_{r_{i}}\backslash\mathcal{O}(\mathscr{P}_{r_{i}})$ do
@@
13:  end for
14:  Merging partitions based on greedy algorithm;
15:  for $i=1$ to $N$ do
16:     Choose $s_{i}\in\mathcal{V}$;
17:     Calculate a set $M_{s_{i}}$ and define $\ell_{s_{i}}=|M_{s_{i}}|$;
18:     ///// /Refer Algorithm 2 to obtain $M_{s_{i}}$
19:     while $j\leq\ell_{s_{i}}$ do
20:        Set the $j$th element of $M_{s_{i}}$ be $J_{s_{i}j}=\{p_{m1},p_{m2},\ldots,p_{m\ell}\}$, where $\ell=|J_{s_{i}j}|$;
21:        if $\bigcup_{k=1}^{\ell}\mathcal{O}_{p_{mk}}=\emptyset$ or $\sum_{k=1}^{\ell}\min\{|\mathcal{O}_{p_{mk}}|,1\}=1$ then
22:           Let $j=j+1$ and go to line $21$;
23:        end if
24:        Set $c_{1}=0$ and $c_{2}=0$;
25:        Calculate $D(\mathscr{P}_{l})$ $\forall l\in\bigcup_{k=1}^{\ell}\mathcal{O}_{p_{mk}}$;
26:        Calculate $D^{*}(\mathscr{P}_{s_{i}})=\left|\bigcup_{k=1}^{\ell}\mathcal{O}_{p_{mk}}\right|$;
27:        Set $c_{1}=1$ if $D^{*}(\mathscr{P}_{s_{i}})-D(\mathscr{P}_{l})\leq D(\mathscr{P}_{s_{i}})-D^{*}(\mathscr{P}_{s_{i}})$ $\forall l\in\bigcup_{k=1}^{\ell}\mathcal{O}_{p_{mk}}$ and $l\neq s_{i}$;
28:        Set $c_{2}=1$ if $\left(D^{*}(\mathscr{P}_{s_{i}})\right)^{2}\leq\sum_{l\in\bigcup_{k=1}^{\ell}\mathcal{O}_{p_{mk}}}D(\mathscr{P}_{l})$;
29:        if $c_{1}c_{2}=1$ then
30:           Let $\mathcal{O}_{p_{m1}}=\bigcup_{k=1}^{\ell}\mathcal{O}_{p_{mk}}$, and $\mathcal{O}_{p_{mk}}=\emptyset$ for $k=2,\ldots,\ell$;
31:        end if
32:        Let $j=j+1$;
33:     end while
34:  end for
@@
(D1) The node constraint conditions can be satisfied. The guarantee of constraint conditions (It is required that the union of all partitions where each agent is located should include all its physical network neighbor agents) is mainly in the first part “establishing partitions” Algorithm 1 because its “Partition initialization” serves the purpose of allowing each node to select partitions to cover its physical network neighboring nodes.
@@
0:  $\mathscr{P}_{s_{i}}$
0:  $M_{s_{i}}$
1:  Initialize $M_{s_{i}}=\emptyset$;
2:  Construct a set of $\mathscr{P}_{s_{i}}$’s subsets, denoted by $\mathscr{P}_{s_{i}}^{c}=\{\mathcal{J}\subset\mathscr{P}_{s_{i}};~{}|\mathcal{J}|\geq 2\}$; Without lossing of generality, define $\mathscr{P}_{s_{i}}^{c}=\{\mathcal{J}_{s_{i}1},\mathcal{J}_{s_{i}2},\ldots,\mathcal{J}_{s_{i}\zeta_{s_{i}}}\}$, where $\zeta_{s_{i}}=2^{|\mathscr{P}_{s_{i}}|}-|\mathscr{P}_{s_{i}}|-1$;
3:  for $k=1$ to $\zeta_{s_{i}}$ do
4:     Calculate $\mathcal{I}(\mathcal{J}_{s_{i}k})=\bigcap_{\mathcal{O}_{p}\in\mathcal{J}_{s_{i}k}}\mathcal{O}_{p}$;
5:     if $|\mathcal{I}(\mathcal{J}_{s_{i}k})|\geq 2$ then
6:        Set $M_{s_{i}}=M_{s_{i}}\cup\{\mathcal{J}_{s_{i}k}\}$;
@@
(D2) The central part of Algorithm 1 adopts the idea of a greedy algorithm. According to (D1), the feasible partition scheme under the constraint conditions has been reached with the first part of Algorithm 1. However, since the first part starts from the node with the smallest degree in the communication network, the node with a large degree (hub node) may have an excessive computational burden. The second part of the algorithm is mainly employed to reduce the computational burden of hub nodes. The developed method allows the hub nodes’ neighbors in the communication network to share the computing burden equally by merging the partitions of hub nodes. Whether it is shared equally depends on:
@@
2) The total dimension of the local observers of the hub node and its neighbors should be reduced compared to that before the merging partition.
Therefore, the merging part will further optimize the partition results obtained by partition initialization.
(D3) Detailed description of “merging partition” part. Suppose a hub node wants to merge partitions $\mathcal{O}_{p_{m1}},\ldots,\mathcal{O}_{p_{m\ell}}$. After the partitions are merged, the local observers of all nodes in the merged partition $\mathcal{O}_{p_{m1}}$ must estimate all the states involved in $\mathcal{O}_{p_{m1}}$. Hence, the dimensions of local observers of all nodes are equal after merging, which can be denoted as $D^{*}(\mathscr{P}_{s_{i}})=|\bigcup_{k=1}^{\ell}\mathcal{O}_{p_{mk}}|$ where $s_{i}$ is the index of the hub node. Therefore, condition 1) in (D2) indicates $D^{*}(\mathscr{P}_{s_{i}})-D(\mathscr{P}_{l})\leq D(\mathscr{P}_{s_{i}})-D^{*}(\mathscr{P}_{s_{i}})$ for $l=1,\ldots,\ell_{s_{i}}$. Furthermore, after merging partitions, there are $D^{*}(\mathscr{P}_{s_{i}})$ nodes in the partition, and each node needs to estimate all states of $D^{*}(\mathscr{P}_{s_{i}})$ nodes. Therefore, condition 2) can be formulated as $\left(D^{*}(\mathscr{P}_{s_{i}})\right)^{2}\leq\sum_{l\in\bigcup_{k=1}^{\ell}\mathcal{O}_{p_{mk}}}D(\mathscr{P}_{l})$. The aforementioned explains the source of step $27$ and step $28$ in Algorithm 1.
Figure 4: Usage example of Algorithm 1: initial partition. The part marked in red is the node with $\ell_{i}\neq 0$, which needs to consider whether to merge partitions.
Figure 5: Usage example of Algorithm 1: Final partition.
@@
First, focus on the first part of the algorithm. The complexity of this part is reflected in the number of calculations of lines $7$ to line $9$. They are the core contents of “establishing partitions” and are calculated with at most $N\cdot\frac{1}{N}\sum_{i=1}^{N}|\mathcal{N}_{i}|$ times. Note that the average degree $\frac{1}{N}\sum_{i=1}^{N}|\mathcal{N}_{i}|$ is scarce relative to the total number $N$ in a large-scale network.
Second, we calculate the time complexity of Algorithm 2. Since Algorithm 2 is the $18$th line of Algorithm 1, its complexity is also a part of the complexity of Algorithm 1. Its core steps include line $4$ and line $6$, and they are calculated with at least $N\cdot\frac{1}{N}\sum_{i=1}^{N}\zeta_{i}$ times. According to the partition selection method, the number of partitions of each node will not exceed its communication network degree. Hence,
@@
Third, core steps (line $25,26,30$) of the last loop (line $15$ to line $34$) of Algorithm 1 should be calculated with $N\cdot\frac{1}{N}\sum_{i=1}^{N}\ell_{s_{i}}$ times. Besides,
@@
The following text explains Algorithm 1 and 2 in more detail with an example. Figure 4 shows the physical and communication network considered in this example, and The detailed steps of partitioning this network are listed in Figure 4 and Figure 5, where the former shows the initial partitioning steps and the latter shows the merging steps. To facilitate readers’ understanding, some key steps will be interpreted.
@@
(I1) We focus on the area containing nodes $1,\ldots,10$. This area is taken as an example to show why partition design should be carried out in the order of $r_{1}\preceq r_{2}\preceq\cdots\preceq r_{N}$. In light of the sorting rule, node $1$ with $\mathcal{N}_{1}=2$ and $\mathcal{C}_{1}=1$ is the first node to be considered. Based on line $7$ to line $10$, we know $\mathcal{O}_{1}=\{1,2,5\}$. Hence, $\mathscr{P}_{1}=\{\mathcal{O}_{1}\}$, $\mathscr{P}_{2}=\{\mathcal{O}_{1}\}$, $\mathscr{P}_{5}=\{\mathcal{O}_{1}\}$. Subsequently, partitions $\mathcal{O}_{2}=\{3,4\}$, $\mathcal{O}_{3}=\{6,7\}$, $\mathcal{O}_{4}=\{5,8,9\}$, $\mathcal{O}_{5}=\{6,10\}$, $\mathcal{O}_{6}=\{10,11\}$ can be selected for nodes $4$, $7$, $8$, and $10$ in order. Then, note that $1,5\in\mathcal{O}(\mathscr{P}_{2})$; thus $\mathcal{O}_{7}=\{2,3,6\}$ is necessary included in $\mathscr{P}_{2}$. Besides, since $\mathcal{N}_{3}$ has been contained in $\mathcal{O}(\mathscr{P}_{3})$, node $3$ needs not to select partitions any more. This reflects the advantages of the proposed node sorting rule: the node with a larger degree in the communication network selects the partition in the later order, which neither affects its partition selection nor interferes with the partition selection of the previous node, but greatly reduces the computational load (most of the required partitions have been selected in advance). The same situation can also be seen at node $6$. However, the unreasonable order of nodes will lead to a large number of partitions with high repetition (See in I2).
@@
(I2) We focus on nodes $42$, $43$, and $44$. Their selection partition order is $42\preceq 43\preceq 44$. Node $42$ selects its partition as $\{40,41,42,43\}$. Hence, only node $44$ remains in $\mathcal{N}_{43}\backslash\mathcal{O}(\mathscr{P}_{43})$. It means $\{43,44\}$ should be added into $\mathscr{P}_{43}$. However, since $41\in\mathcal{N}_{44}$, $\mathscr{P}_{44}$ must include a partition $\{41,43,44\}$. It leads to a new partition to be added to node $43$, which has completed the partition selection before that. More unreasonably, two highly repetition partitions $\{41,43,44\},\{43,44\}$ exist at the same time in $\mathscr{P}_{43}$ and $\mathscr{P}_{44}$. The reason for this phenomenon is that $|\mathcal{N}_{44}\backslash\mathcal{O}(\mathscr{P}_{44})|>|\mathcal{N}_{43}\backslash\mathcal{O}(\mathscr{P}_{43})|$ after the node $42$ selects its partitions. The above analysis shows that we have the advantage of allowing the node with a smaller communication network degree and larger physical network degree to select partitions preferentially. Otherwise, highly similar partitions like $\{41,43,44\}$ and $\{43,44\}$ will appear in large numbers. In addition, the above analysis also shows that a more reasonable sorting rule should be: node $i$ will preferentially select the partition if $|\mathcal{N}_{i}\backslash\mathcal{O}(\mathscr{P}_{i})|>|\mathcal{N}_{j}\backslash\mathcal{O}(\mathscr{P}_{j})|$ when $|\mathcal{C}_{j}|=|\mathcal{C}_{i}|$. However, this sort rule means that one has to sort the nodes again after a node selects its partition, which is adverse to the implementation of the algorithm. Therefore, we do not improve the algorithm from the perspective of the remaining physical network neighbor nodes but use the merging strategy to make up for this problem.
@@
(I3) Figure 5 displays the steps of merging partitions. Note that the actual calculation is not complex, although the complexity of the partition merging part is $O(N^{2})$. It is because after the partitions of some nodes are merged, many other nodes that initially needed to merge partitions no longer need to perform the merging steps. For example, in Figure 4, all nodes $41,43,44$ need to merge partitions. However, $M_{s_{i}}$ (defined by Algorithm 2) with respect to node $43$ and $44$ are both empty after $41$ merging its partitions, and thus they need not perform the merging steps anymore.

### v4

@@ III-A Cover sets selecting
Algorithm 1 Coverage solving algorithm of distributed observer for large-scale systems
[…]
0: A cover set of distributed observer
1:Establishing cover sets based on greedy algorithm;
2: Initialize $p=1$ and $\mathcal{O}_{p}=\emptyset$; initialize $\mathscr{P}_{i}=\emptyset$ $\forall i\in\mathcal{V}$; initialize $\mathcal{V}_{r}=\mathcal{V}$ and $\mathcal{V}_{r}=\mathcal{V}$;
3:while $|\mathcal{V}_{r}|\neq 0$ do
4:  Choose $i\in\mathcal{V}_{r}$ with minimum $r_{i}$;
5:  if $\mathcal{N}_{i}\backslash\mathcal{O}(\mathscr{P}_{r_{i}})\neq\emptyset$ then
6:   for $j\in\mathcal{N}_{i}\backslash\mathcal{O}(\mathscr{P}_{r_{i}})$ do
@@
[…]
14:end while
15:Merging cover sets based on greedy algorithm;
16:while $|\mathcal{V}_{s}|\neq 0$ do
17:  Choose $i\in\mathcal{V}_{s}$ with maximum $s_{i}$;
18:  Calculate a set $M_{i}$ and define $\ell_{i}=|M_{i}|$;
19:  ////Refer Algorithm 1 to obtain $M_{i}$
20:  while $j\leq\ell_{i}$ do
21:   Set the $j$th element of $M_{i}$ be $J_{ij}=\{p_{m1},p_{m2},\ldots,p_{m\ell}\}$, where $\ell=|J_{ij}|$;
22:   if $\bigcup_{k=1}^{\ell}\mathcal{O}_{p_{mk}}=\emptyset$ or $\sum_{k=1}^{\ell}\min\{|\mathcal{O}_{p_{mk}}|,1\}=1$ then
23:    Let $j=j+1$ and go to line 2121;
24:   end if
25:   Set $c_{1}=0$ and $c_{2}=0$;
26:   Calculate $D(\mathscr{P}_{l})$ $\forall l\in\bigcup_{k=1}^{\ell}\mathcal{O}_{p_{mk}}$;
27:   Calculate $D^{*}(\mathscr{P}_{s_{i}})=\left|\bigcup_{k=1}^{\ell}\mathcal{O}_{p_{mk}}\right|$;
28:   Set $c_{1}=1$ if $D^{*}(\mathscr{P}_{s_{i}})-D(\mathscr{P}_{l})\leq D(\mathscr{P}_{s_{i}})-D^{*}(\mathscr{P}_{s_{i}})$ $\forall l\in\bigcup_{k=1}^{\ell}\mathcal{O}_{p_{mk}}$ and $l\neq i$;
29:   Set $c_{2}=1$ if $\left(D^{*}(\mathscr{P}_{s_{i}})\right)^{2}\leq\sum_{l\in\bigcup_{k=1}^{\ell}\mathcal{O}_{p_{mk}}}D(\mathscr{P}_{l})$;
30:   if $c_{1}c_{2}=1$ then
31:    Let $\mathcal{O}_{p_{m1}}=\bigcup_{k=1}^{\ell}\mathcal{O}_{p_{mk}}$, and $\mathcal{O}_{p_{mk}}=\emptyset$ for $k=2,\ldots,\ell$;
32:   end if
33:   Let $j=j+1$;
34:  end while
[…]
36:end while
Now, one can see Algorithm 1 for the developed coverage solving method. To aid the reader’s understanding, the following is a brief description of the algorithm.
Step 1: Sort $r_{i},~i=1,\ldots,N$ based on symbol $\preceq$.
Step 2: Each node traverses its physical network neighbors. If a neighbor is already in $\mathcal{O}(\mathscr{P}_{i})$, no action is taken. If the neighbor is not in $\mathcal{O}(\mathscr{P}_{i})$, all nodes on the shortest path from the current node to that neighbor are included in a new subset $\mathcal{O}_{p}$ (subsequently, any other newly added physical network neighbors and the nodes on their shortest paths will also be placed in this same $\mathcal{O}_{p}$; that is, each node creates only one new subset).
Step 3: Sort $s_{i},~i=1,\ldots,N$ based on symbol $\succeq$.
Step 4: Starting from the maximum $s_{i}$. Each node considers whether to merge several cover sets according to two criteria (Line 27–29 in Algorithm 1 and their details are shown in (D2)) in sequence (In communication networks, nodes with larger degrees are more likely to be forced to join many redundant cover sets. Therefore, priority is given to nodes with larger degrees to determine whether to remove some redundant cover sets through merging).
Note that the coverage solving algorithms described in Algorithm 1 and Algorithm 2 actually solve a multi-objective optimization problem because each node has its own desired cover sets, but the desired optimal coverage among different nodes are conflicting. In the complete version of this article, we prove that the coverage obtained through Algorithm 1 and Algorithm 2 are Pareto solutions to this multi-objective optimization problem.
Propsition 1
Coverage obtained from Algorithm 1 and Algorithm 2 are already located on the Pareto frontier.
Proof:
Given a group of cover sets ($\mathcal{O}_{1},\ldots,\mathcal{O}_{m}$) obtained by Algorithm 1 and Algorithm 2. We first prove three propositions.
1) Add any nodes in arbitrary $\mathcal{O}_{i}=\{i_{1},\ldots,i_{r}\},~i\in\{1,2,\ldots,m\}$ would increase the dimensions of observers on at least one node. Since $\mathcal{O}_{i}$ is obtained by at least one node (denote as $i_{1}$) using a greedy algorithm, node $i_{1}$ makes the optimal choice when selecting $\mathcal{O}_{i}$ based on the current situation. Now, let’s discuss two scenarios. First, $i_{1}$ has yet to be included in any cover set before selecting $\mathcal{O}_{i}$. In this case, $\mathcal{O}_{i}$ is the first cover set for $i_{1}$, so it is the optimal choice for $i_{1}$ at that time. In this scenario, adding any new node to this cover set will increase the dimension of the observer on $i_{1}$. Second, before selecting $\mathcal{O}_{i}$, $i_{1}$ has already been included in another cover set (denote as $\mathcal{O}_{j}$) by another node (let’s call it $i_{2}$). In this case, if a new node $i_{0}$ is added into $\mathcal{O}_{i}$, then the dimension of observers on node $i_{1}$ will increase owing to the joining of $i_{0}$, unless $\mathcal{O}_{j}$ becomes a subset of $\mathcal{O}_{i}\cup\{i_{0}\}$ and node $i$ chooses to merge $\mathcal{O}_{i}$ and $\mathcal{O}_{j}$. However, $\mathcal{O}_{j}$ is the optimal choice made by $i_{2}$ based on the greedy algorithm. Therefore, any changes made by $i_{1}$ to $\mathcal{O}_{j}$ will inevitably result in an increase in the observer dimension on $i_{2}$. This proposition is proven.
[…]
Based on the above three propositions, we will prove that the coverage obtained by Algorithm 1 and Algorithm 2 are already Pareto optimal solutions. To this end, it is sufficient to show that reducing the value of $D(\mathscr{P}_{j})$ for any node (e.g., $j$) will result in an increase in $D(\mathscr{P}_{k})$ for at least one node (e.g., $k$). We will discuss this with two cases.
First, let’s assume that $\mathscr{P}_{j}$ contains multiple cover sets ($\mathcal{O}_{j1}$, $\mathcal{O}_{j2}$, $\ldots$). In order to reduce $D(\mathscr{P}_{j})$, node $j$ can either remove a node from a certain set (e.g., deleting node $k$ from $\mathcal{O}_{j1}$) or merge multiple cover sets (e.g., merging $\mathcal{O}_{j1}$ and $\mathcal{O}_{j2}$). Based on the previous propositions, there exists at least one node $i_{1}$, for which $\mathcal{O}_{j1}$ is optimal. Therefore, any operation of reducing or merging $\mathcal{O}_{j1}$ will inevitably lead to an increase in $D(\mathscr{P}_{i_{1}})$. Hence, in this case, optimizing the observer dimensions of node $j$ will inevitably result in at least one other node suffering a loss in benefit or resulting in an invalid coverage.
Second, let’s assume that $\mathscr{P}_{j}$ contains only one cover set. In this case, the only way for node $j$ to reduce $D(\mathscr{P}_{j})$ is by removing some nodes from that cover set. However, according to Proposition 2), this approach will obviously result in at least one other node suffering a loss in benefit.
In conclusion, the coverage obtained from Algorithm 1 and Algorithm 2 are already located on the Pareto frontier. ∎
We have the following discussions regarding Algorithm 1 and 2.
@@
(D1) The node constraint conditions can be satisfied. The guarantee of constraint conditions (It is required that the union of all cover sets where each agent is located should include all its physical network neighbor agents) is mainly in the first part “establishing cover sets” Algorithm 1 because its “cover sets initialization” serves the purpose of allowing each node to select sets to cover its physical network neighboring nodes.
@@
[…]
1: Initialize $M_{i}=\emptyset$;
2: Construct a set of $\mathscr{P}_{i}$’s subsets, denoted by $\mathscr{P}_{i}^{c}=\{\mathcal{J}\subset\mathscr{P}_{i};~|\mathcal{J}|\geq 2\}$; Without lossing of generality, define $\mathscr{P}_{i}^{c}=\{\mathcal{J}_{i1},\mathcal{J}_{i2},\ldots,\mathcal{J}_{i\zeta_{i}}\}$, where $\zeta_{i}=2^{|\mathscr{P}_{i}|}-|\mathscr{P}_{i}|-1$;
3:for $k=1$ to $\zeta_{i}$ do
4:  Calculate $\mathcal{I}(\mathcal{J}_{ik})=\bigcap_{\mathcal{O}_{p}\in\mathcal{J}_{ik}}\mathcal{O}_{p}$;
5:  if $|\mathcal{I}(\mathcal{J}_{ik})|\geq 2$ then
6:   Set $M_{i}=M_{i}\cup\{\mathcal{J}_{ik}\}$;
@@
(D2) The central part of Algorithm 1 adopts the idea of a greedy algorithm. According to (D1), the feasible coverage under the constraint conditions has been reached with the first part of Algorithm 1. However, since the first part starts from the node with the smallest degree in the communication network, the node with a large degree (hub node) may have an excessive computational burden. The second part of the algorithm is mainly employed to reduce the computational burden of hub nodes. The developed method allows the hub nodes’ neighbors in the communication network to share the computing burden equally by merging the cover sets of hub nodes. Whether it is shared equally depends on:
@@
2) The total dimension of the local observers of the hub node and its neighbors should be reduced compared to that before merging cover sets.
Therefore, the merging part will further optimize the coverage obtained by cover sets initialization.
(D3) Description of the “Merging Cover Set” part. Suppose a hub node wants to merge cover sets $\mathcal{O}_{p_{m1}},\ldots,\mathcal{O}_{p_{m\ell}}$. After the sets are merged, the local observers of all nodes in the merged set $\mathcal{O}_{p_{m1}}$ must estimate all the states involved in $\mathcal{O}_{p_{m1}}$. Hence, the dimensions of local observers of all nodes are equal after merging, which can be denoted as $D^{*}(\mathscr{P}_{s_{i}})=|\bigcup_{k=1}^{\ell}\mathcal{O}_{p_{mk}}|$ where $s_{i}$ is the index of the hub node. Therefore, condition 1) in (D2) indicates $D^{*}(\mathscr{P}_{s_{i}})-D(\mathscr{P}_{l})\leq D(\mathscr{P}_{s_{i}})-D^{*}(\mathscr{P}_{s_{i}})$ for $l=1,\ldots,\ell_{s_{i}}$. Furthermore, after merging cover sets, there are $D^{*}(\mathscr{P}_{s_{i}})$ nodes in the set, and each node needs to estimate all states of $D^{*}(\mathscr{P}_{s_{i}})$ nodes. Therefore, condition 2) can be formulated as $\left(D^{*}(\mathscr{P}_{s_{i}})\right)^{2}\leq\sum_{l\in\bigcup_{k=1}^{\ell}\mathcal{O}_{p_{mk}}}D(\mathscr{P}_{l})$. The aforementioned explains the source of step $27$ and step $28$ in Algorithm 1.
Figure 4: Usage example of Algorithm 1: initial cover set. The part marked in red is the node with $\ell_{i}\neq 0$, which needs to consider whether to merge covers sets.
Figure 5: Usage example of Algorithm 1: Final coverage.
@@
First, focus on the first part of the algorithm. The complexity of this part is reflected in the number of calculations of lines $7$ to line $8$. They are the core contents of “establishing coverage” and are calculated with at most $N\cdot\frac{1}{N}\sum_{i=1}^{N}|\mathcal{N}_{i}|$ times. Note that the average degree $\frac{1}{N}\sum_{i=1}^{N}|\mathcal{N}_{i}|$ is scarce relative to the total number $N$ in a large-scale network.
Second, we calculate the time complexity of Algorithm 2. Since Algorithm 2 is the $19$th line of Algorithm 1, its complexity is also a part of the complexity of Algorithm 1. Algorithm 2’s core steps include line $4$ and line $6$, and they are calculated with at least $N\cdot\frac{1}{N}\sum_{i=1}^{N}\zeta_{i}$ times. According to the coverage solving method, the number of cover sets of each node will not exceed its communication network degree. Hence,
@@
Third, core steps (line $26,27,31$) of the last loop (line $16$ to line $36$) of Algorithm 1 should be calculated with $N\cdot\frac{1}{N}\sum_{i=1}^{N}\ell_{s_{i}}$ times. Additionally,
@@
The following text explains Algorithm 1 and 2 in more detail with an example. Figure 4 shows the physical and communication network considered in this example, and The detailed steps of find coverage are listed in Figure 4 and Figure 5, where the former shows the initial cover sets steps and the latter shows the merging steps. To facilitate readers’ understanding, some key steps will be interpreted.
@@
(I1) We focus on the area containing nodes $1,\ldots,10$. This area is taken as an example to show why cover set selecting should be carried out in the order of $r_{1}\preceq r_{2}\preceq\cdots\preceq r_{N}$. In light of the sorting rule, node $1$ with $\mathcal{N}_{1}=2$ and $\mathcal{C}_{1}=1$ is the first node to be considered. Based on line $7$ to line $10$, we know $\mathcal{O}_{1}=\{1,2,5\}$. Hence, $\mathscr{P}_{1}=\{\mathcal{O}_{1}\}$, $\mathscr{P}_{2}=\{\mathcal{O}_{1}\}$, $\mathscr{P}_{5}=\{\mathcal{O}_{1}\}$. Subsequently, sets $\mathcal{O}_{2}=\{3,4\}$, $\mathcal{O}_{3}=\{6,7\}$, $\mathcal{O}_{4}=\{5,8,9\}$, $\mathcal{O}_{5}=\{6,10\}$, $\mathcal{O}_{6}=\{10,11\}$ can be selected for nodes $4$, 77, 88, and $10$ in order. Then, note that $1,5\in\mathcal{O}(\mathscr{P}_{2})$; thus $\mathcal{O}_{7}=\{2,3,6\}$ is necessary included in $\mathscr{P}_{2}$. Besides, since $\mathcal{N}_{3}$ has been contained in $\mathcal{O}(\mathscr{P}_{3})$, node $3$ needs not to select sets any more. This reflects the advantages of the proposed node sorting rule: the node with a larger degree in the communication network selects the set in the later order, which neither affects its cover set selection nor interferes with the cover set selection of the previous node, but greatly reduces the computational load (most of the required cover sets have been selected in advance). The same situation can also be seen at node $6$. However, the unreasonable order of nodes will lead to a large number of cover sets with high repetition (See in I2).
@@
(I2) We focus on nodes $42$, $43$, and $44$. Their selection order is $42\preceq 43\preceq 44$. Node $42$ selects its cover sets as $\{40,41,42,43\}$. Hence, only node $44$ remains in $\mathcal{N}_{43}\backslash\mathcal{O}(\mathscr{P}_{43})$. It means $\{43,44\}$ should be added into $\mathscr{P}_{43}$. However, since $41\in\mathcal{N}_{44}$, $\mathscr{P}_{44}$ must include a set $\{41,43,44\}$. It leads to a new set to be added to node $43$, which has completed the cover set selection before that. More unreasonably, two highly repetition cover sets $\{41,43,44\},\{43,44\}$ exist at the same time in $\mathscr{P}_{43}$ and $\mathscr{P}_{44}$. The reason for this phenomenon is that $|\mathcal{N}_{44}\backslash\mathcal{O}(\mathscr{P}_{44})|>|\mathcal{N}_{43}\backslash\mathcal{O}(\mathscr{P}_{43})|$ after the node $42$ selects its cover sets. The above analysis shows that we have the advantage of allowing the node with a smaller communication network degree and larger physical network degree to select cover sets preferentially. Otherwise, highly similar cover sets like $\{41,43,44\}$ and $\{43,44\}$ will appear in large numbers. In addition, the above analysis also shows that a more reasonable sorting rule should be: node $i$ will preferentially select the cover set if $|\mathcal{N}_{i}\backslash\mathcal{O}(\mathscr{P}_{i})|>|\mathcal{N}_{j}\backslash\mathcal{O}(\mathscr{P}_{j})|$ when $|\mathcal{C}_{j}|=|\mathcal{C}_{i}|$. However, this sort rule means that one has to sort the nodes again after a node selects its cover sets, which is adverse to the implementation of the algorithm. Therefore, we do not improve the algorithm from the perspective of the remaining physical network neighbor nodes but use the merging strategy to make up for this problem.
@@
(I3) Figure 5 displays the steps of merging cover sets. Note that the actual calculation is not complex, although the complexity of the merging part is $O(N^{2})$. It is because after the cover sets of some nodes are merged, many other nodes that initially needed to merge cover sets no longer need to perform the merging steps. For example, in Figure 4, all nodes $41,43,44$ need to merge their cover sets. However, $M_{s_{i}}$ (defined by Algorithm 2) with respect to node $43$ and $44$ are both empty after $41$ merging its sets, and thus they need not perform the merging steps anymore.